OpenStudy (anonymous):

.What is 80 ÷ (-50)?

OpenStudy (bahrom7893):

-80/50 = -8/5

OpenStudy (anonymous):

-1.75

OpenStudy (anonymous):

wait -1.6

OpenStudy (anonymous):

what about as a fraction ?

OpenStudy (anonymous):

bahroms answer

OpenStudy (anonymous):

easy way to think about it is: -80/-50 = (-8*10)/(-5/*10)....the 10's will cancel here =-8/5

OpenStudy (anonymous):

-8/5